OpenStudy (anonymous):

Lim f(x) = (x)/ (x^2 - x). It has discontinuities at; a. nonremovable at x= 0 b. nonremovable at x = 0; removable at x = 1 c. removable at x = 0 d. removable at x=0 ; nonremovable at x = 1 e. the graoh is continuous for all real numbers

OpenStudy (anonymous):

see if you can remove a discontinuity by factoring and cancelling, because clearly this is undefined if \(x=0\) or \(x=1\)

OpenStudy (anonymous):

oops, somehow my post got deleted, but i hope the answer was clear. you get \[f(x)=\frac{x}{x^2-x}=\frac{x}{x(x-1)}=\frac{1}{x-1} \text{ if } x\neq 0\]

OpenStudy (anonymous):

so zero is a removable discontinuity but 1 is not removable